Differentiate the following function:g(x)= x3+(1/x3)





1
Expert's answer
2021-02-16T13:49:28-0500

Solution.

"g(x)= x^3+\frac{1}{x^3}"

"g'(x)=3x^2-3x^{-4}=3x^2-\frac{3}{x^4}.
\newline"

Derivatives in this expression are calculated using the formula:


"(x^n)'=n\cdot x^{n-1}."

Answer. "g'(x)=3x^2-\frac{3}{x^4}."